OpenStudy (legends):

Solve the system of equations. 5x+4=9 3x+2y=4

OpenStudy (anonymous):

are you solving for both x?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

is this a trick question?

OpenStudy (legends):

no

OpenStudy (misty1212):

it is really \[5x+4=9\\ 3x+2y=4\] or is it \[5+4y=9\\ 3x+2y=4\]?

OpenStudy (anonymous):

5x + 4 = 9 3x + 2y = 4 Are you missing 'y' in the first equation? @Legends Is there supposed to be a 'y' after '4'?

OpenStudy (legends):

Sorry guys yea its 5x+y=9

OpenStudy (legends):

@iGreen. Yea its 5x + 4 = 9 3x + 2y = 4

OpenStudy (anonymous):

Oh, okay.. 5x + y = 9 3x + 2y = 4 This is it?

OpenStudy (legends):

yea

OpenStudy (anonymous):

We can start by subtract 5x to both sides in the first equation: y = -5x + 9 Now we can substitute '-5x + 9' for 'y' in the 2nd equation: 3x + 2y = 4 3x + 2(-5x + 9) = 4 Distribute 2 into the parenthesis: 3x + -10x + 18 = 4 Combine: -7x + 18 = 4 Subtract 18 to both sides, what's 4 - 18?

OpenStudy (legends):

-14

OpenStudy (anonymous):

Yes, so we have: -7x = -14 Divide -7 to both sides, what's -14 / -7?

OpenStudy (legends):

2

OpenStudy (legends):

So is the answer(2,-1)

OpenStudy (legends):

@iGreen.

OpenStudy (anonymous):

Hold on.

OpenStudy (anonymous):

The x-value is 2..

OpenStudy (anonymous):

Yep, you got it.

OpenStudy (legends):

thanks

OpenStudy (anonymous):

We plug 2 back into any of the two equations for 'x' and we get -1 for y..